OpenStudy (anonymous):

Evaluate- 2*2^3*2^-5*2^-1

OpenStudy (anonymous):

2^(1+3-5-1) =2^-2 =1/2^2 =1/4

OpenStudy (anonymous):

Thank you^

OpenStudy (anonymous):

yw